# Assortment Optimization Under History-Dependent Effects

### v2

@@ 4.4 A Strong Formulation for Finding Two-Cyclic Policies
Note that this reformulation reveals a connection between (Cycle) and the assortment problem studied in Caro et al. (2014) in which a retailer must decide, in advance, the release date of each product in a given collection over a selling season. More specifically, (2-Cyclic) is exactly the case where a retailer faces two planning horizons and sells products with a single period life-cycle, and, in each period, customers make a purchase decision on offered products according to the MNL model.
@@
Next, we introduce variables for building a strong formulation of (2-Cyclic). For each period $t=1,2$, we use $\rho^{t}$ and $\gamma^{t}_{i}$ to represent the no-purchase probability and choice probability of product $i$, respectively
$\rho^{t}=\frac{1}{1+\sum_{j\in[N]}u_{j}x^{t}_{j}}\quad\text{ and }\quad\gamma_{i}^{t}=\frac{x^{t}_{i}}{1+\sum_{j\in[N]}u_{j}x^{t}_{j}}.\qquad\operatorname{for}i\in[N]$ (CC)
@@
Note that this is known as the Charnes-Cooper transformation (Charnes and Cooper 1962). For each period $t=1,2$, it maps all possible assortments $\boldsymbol{x}^{t}\in\{0,1\}^{N}$ into all possible choice probabilities $(\rho^{t},u_{1}\gamma^{t}_{1},\ldots,u_{N}\gamma^{t}_{N})$, where $(\rho^{t},\boldsymbol{\gamma}^{t})$ belongs to
$\rho^{t}+\sum_{i\in[N]}u_{i}\gamma^{t}_{i}=1\quad\text{ and }\quad\gamma^{t}_{i}\in\{0,\rho^{t}\}\,\operatorname{for}i\in[N].$ (10)
Using these relations, we are ready to present a new formulation of (2-Cyclic):
$\max\quad$ $\sum_{t=1,2}\frac{1}{2}\sum_{i\in[N]}r_{i}u_{i}\gamma_{i}^{t}$
$\operatorname{s.t.}\quad$ $x^{1}_{i}+x^{2}_{i}\leq 1\text{ and }\boldsymbol{x}^{t}\in\mathcal{X}\{0,1\}^{N}$ $\operatorname{for}i\in[N]\text{ and }t=1,2$ (11a)
$\rho^{t}+\sum_{i\in[N]}u_{i}\gamma^{t}_{i}=1\text{ and }0\leq\gamma^{t}_{i}\leq\rho^{t}\quad$ $\operatorname{for}i\in[N]\text{ and }t=1,2$ (11b)
$x^{t}_{i}=\gamma^{t}_{i}+\sum_{j\in[N]}u_{j}\Gamma^{t}_{ij}$ $\operatorname{for}i\in[N]\text{ and }t=1,2$ (11c)
$\max\{0,\gamma^{t}_{i}+\gamma^{t}_{j}-\rho^{t}\}\leq\Gamma^{t}_{ij}\leq\min\{\gamma^{t}_{i},\gamma^{t}_{j}\}\quad$ $\operatorname{for}i\in[N]\,,j\in[N]\setminus\{i\},\text{ and }t=1,2$ (11d)
$\Gamma^{t}_{ii}=\gamma^{t}_{i}$ $\operatorname{for}i\in[N]\text{ and }t=1,2.$ (11e)
@@
The constraint (11b) is obtained from relaxing the constraints from (10). The constraint (11c) is derived from the following equality, with the variable $\gamma^{t}_{i}$ replacing the first term and a new variable $\Gamma^{t}_{ij}$ replacing each fractional term in the summation,
$x^{t}_{i}=\frac{x^{t}_{i}}{1+\sum_{k\in[N]}u_{k}x^{t}_{k}}+\sum_{j\in[N]}u_{j}\frac{x_{i}^{t}x_{j}^{t}}{1+\sum_{k\in[N]}u_{k}x^{t}_{k}}\qquad\operatorname{for}\,t=1,2.$
To obtain the constraint (11d), we use the no-purchasee variable $\rho^{t}$ to scale the following McCormick envelope of $x_{i}^{t}x_{j}^{t}$ (McCormick 1976):
$\max\{0,x^{t}_{i}+x^{t}_{j}-1\}\leq x^{t}_{i}x^{t}_{j}\leq\min\{x_{i}^{t},x_{j}^{t}\}\qquad\operatorname{for}j\in[N].$
@@
The last constraint (11e) is obtained by observing that the relation $x^{t}_{i}x^{t}_{i}=x^{t}_{i}$ holds for binary variables. Unlike the formulation (Cycle-Conic), this formulation does not rely on bounds of the no-purchasing probability variables. More specifically, in the derivation of (Cycle-Conic), we use $\rho^{t}x^{t}_{i}$ as a nonlinear scaling factor, demanding a linearization using bounds of $\rho^{t}$. In the development of linear inequalities in (11), we avoid using nonlinear scaling factor and, instead, use the no-purchasing probability variables, $(\rho^{1},\rho^{2})$, as scaling variables to derive (11d).
@@
The following result shows this bounds-free formulation is theoretically tighter than the formulation (Cycle-Conic). One of the most important properties of a MIP formulation is the strength of its natural continuous relaxation that is obtained by ignoring the integrality constraints. This is important because a tighter continuous relaxation often indicates a faster convergence of the branch-and-bound algorithm, on which most commercial MIP solvers are built (Vielma 2015). Since (HAP) has a maximization objective, a tighter formulation has a smaller continuous relaxation objective value. Let $Z_{\textsc{Bounds-free}}$ (resp. $Z_{\textsc{Cycle-Conic}}$) be the optimal objective value of the natural continuous relaxation of formulation (11) (resp. (Cycle-Conic)). The two objective values have the following relationship.
@@
For $M=1$, $Z_{\textsc{Bounds-free}}\leq Z_{\textsc{Cycle-Conic}}$ if the cross-period constraint is non-overlapping.
@@ 5 Case Study
This section presents a case study using transaction data from a corporation cafeteria. We first estimate the history-dependent effect using the historical data and then use the results to design assortments. Our estimations indicate that customers have a significant but short-term satiation effect. Furthermore, our findings suggest that including history-dependent effects in choice models will lead to greater assortment varieties and higher revenue.
@@ 5.1 Data Description and Estimation
We cooperated with a Chinese company that provides catering services to over $4,000$ corporation cafeterias and were given access to a dataset for one cafeteria that mainly serves workday luncheons for customers who work nearby. [^6] The cafeteria is closed on holidays and weekends. The dataset includes $27,144$ transactions for $53$ main dishes, covering $110$ workdays from July to December, 2021. The data for each transaction include the time, dish, and quantity, as well as a card ID if the dish was purchased with a membership card. The dataset also includes information about the price, main ingredient, and flavor of each dish. In what follows, we use the terms dish and product and also the terms menu and assortment interchangeably.
Our dataset indicates that 84.8% of the transactions came from 418 members, which generated 84.5% of the total revenue. Thus, we focus on the members’ transactions because they are frequent visitors and generate most of the revenue. Our dataset also reveals that the probability of a member dining in the cafeteria more than once for two, three, or four consecutive days was 72.5%, 87.7%, and 91.2%, respectively. This suggests that customers’ dish selections could be influenced by their past dining experiences.
@@
$U_{i}^{t}=\boldsymbol{\theta}^{T}X_{i}^{t}+\theta_{r}r_{i}+\sum_{m\in[M]}x_{i}^{t-m}\beta^{m}+\xi_{i}^{t},$
@@
where $X_{i}^{t}$ denotes covariates that include the main ingredient and flavor of product $i$, product fixed effects, and month fixed effects; $r_{i}\in[0,1]$ is a normalized price; $\boldsymbol{\theta}$ and $\theta_{r}$ are coefficients of $X_{i}^{t}$ and price, respectively; $\boldsymbol{\theta}^{T}X_{i}^{t}+\theta_{r}r_{i}$ is the base utility $\beta_{i}^{0}$ defined in Section 2; and the $\xi_{i}^{t}$s are i.i.d. standard Gumbel random variables. The history-dependent effect is captured by the term $\sum_{m\in[M]}x_{i}^{t-m}\beta^{m}$, in which $M$ is the memory length. Our model is a standard MNL model if $M=0$. Note that we set the same history-dependent coefficient for all products because not all products have consecutive offering records and therefore we cannot estimate the product-wise history-dependent effect. We do not include the cross-product history-dependent effect because the estimation result shows that the cross-product effect is approximately one-tenth of the direct effect and is significant only when $M=1$. This is in line with numerous empirical findings that the direct-product effect is much more significant than the cross-product effect (e.g., Vilcassim et al. 1999, Dubé and Manchanda 2005, Mehta 2007). Therefore, we focus on the direct history-dependent effect.
@@
We utilize members’ transactional data to calibrate the MNL choice model with the utility function presented above. We split the data into two sets: transaction data from July to September and transaction data from October to December. We first use five-fold cross-validation on the first data set to estimate the above utility function, and then we optimize weekly menus based on the estimators. Next, we conduct the five-fold cross-validation again, this time using the second data set, to select the value of $M$ with the best out-of-sample performance, and we set this as the “ground truth” model for evaluating revenue, which is a common approach in policy evaluation (Cao et al. 2022). To address the fact that our dataset lacks no-purchase records, we employ the expectation-maximization (EM) approach, which is a widely used technique to deal with censored data (Talluri and Van Ryzin 2006, van Ryzin and Vulcano 2017, Şimşek and Topaloglu 2018), to calibrate our model.
@@
Table 3 summarizes the estimation and five-fold validation results for $M=\{0,1,2,3,4\}$, using the third quarter data. Column $\theta_{r}$ presents price coefficient estimators, and columns $\beta^{1}$ to $\beta^{4}$ give history-dependent effects. The columns labelled $\chi^{2}$ and Log-likelihood provide the average out-of-sample value in the five-fold cross-validation (normalized by sample size). The percentage values in parentheses indicate the relative improvement of each model with a history-dependent effect compared to the standard MNL model.
Table 3: Estimators and cross-validation results
Estimators Out-of-sample Performance
$M$ $\theta_{r}$ $\beta^{1}$ $\beta^{2}$ $\beta^{3}$ $\beta^{4}$ $\chi^{2}$ Log-likelihood
0 $-2.493^{***}$ 0.0916 -1.7985
1 $-2.468^{***}$ $-0.480^{***}$ 0.0900 ($1.75\%$) -1.7966 ($0.10\%$)
(0.082) (0.041)
2 $-2.444^{***}$ $-0.494^{***}$ $-0.322^{***}$ 0.0890 ($2.84\%$) -1.7953 ($0.18\%$)
(0.082) (0.041) (0.040)
3 $-2.440^{***}$ $-0.494^{***}$ $-0.307^{***}$ $-0.225^{***}$ 0.0896 ($2.18\%$) -1.7951 ($0.19\%$)
(0.082) (0.041) (0.041) (0.052)
4 $-2.433^{***}$ $-0.499^{***}$ $-0.307^{***}$ $-0.222^{***}$ $-0.106^{***}$ 0.0905 ($1.20\%$) -1.7955 ($0.17\%$)
(0.082) (0.041) (0.041) (0.052) (0.045)
@@
  • Note: The numbers in parentheses under the estimators report the standard errors of the estimations. “-” indicates that a model does not include the corresponding predictor. ${***}p<0.001$. The columns “$\chi^{2}$” and “Log-likelihood” show the average value of the five-fold cross-validation (normalized by sample size) followed by a percentage in parentheses, which is the relative improvement of each model with a history-dependent effect compared to the standard MNL model.
@@
Table 3 shows significant negative history-dependent effects, and the scale of the effects decrease as the time interval between the current and historical offerings increases. The estimation results are consistent for different values of $M$. For instance, $\beta^{1}$ is consistently about $-0.49$ for all $M$. This suggests that the attraction value of a product will shrink by about $\exp(-0.49)=61\%$ if it was offered yesterday. In other words, if a product is sold on two consecutive days, we need to lower its price by at least RMB $(0.49/2.493)\times 15\approx 2.95$ to compensate for the negative history-dependent effect. [^7] We normalize prices by the maximum price RMB $15$ in the estimation. These results demonstrate the importance of considering history-dependent effects in assortment planning.
@@
The average out-of-sample validation results are reported in the last two columns of Table 3. They suggest that history-dependent effects can improve prediction power (i.e., lower $\chi^{2}$s and higher log-likelihoods). When $M=2$, the model presents the lowest $\chi^{2}$ and the second highest log-likelihood. This result suggests that a small memory length ($M=2$) can capture the history-dependent effect in customers’ choice.

### v3

@@ 4.4 A Strong Formulation for Finding Non-overlapping $(M+1)$-Cyclic Policies
Note that this reformulation reveals a connection between (Cycle) and the assortment problem studied in Caro et al. (2014) in which a retailer must decide, in advance, the release date of each product in a given collection over a selling season. More specifically, when $M=1$, ((M+1)-Cyclic) is exactly the case where a retailer faces two planning horizons and sells products with a single period life-cycle, and, in each period, customers make a purchase decision on offered products according to the MNL model.
@@
Next, we introduce variables for building a strong formulation of ((M+1)-Cyclic). For each period $t\in[M+1]$, we use $\rho^{t}$ and $u_{i}\cdot\gamma^{t}_{i}$ to represent the no-purchase probability and choice probability of product $i$, respectively,
$\rho^{t}=\frac{1}{1+\sum_{j\in[N]}u_{j}x^{t}_{j}}\quad\text{ and }\quad\gamma_{i}^{t}=\frac{x^{t}_{i}}{1+\sum_{j\in[N]}u_{j}x^{t}_{j}}\qquad\text{ for }i\in[N].$ (CC)
@@
Note that this is known as the Charnes-Cooper transformation (Charnes and Cooper 1962). For each period $t\in[M+1]$, it maps all possible assortments $\boldsymbol{x}^{t}\in\{0,1\}^{N}$ into all possible choice probabilities $(\rho^{t},u_{1}\gamma^{t}_{1},\ldots,u_{N}\gamma^{t}_{N})$, where $(\rho^{t},\boldsymbol{\gamma}^{t})$ belongs to
$\rho^{t}+\sum_{j\in[N]}u_{j}\gamma^{t}_{j}=1\quad\text{ and }\quad\gamma^{t}_{j}\in\{0,\rho^{t}\}\,\operatorname{for}j\in[N].$ (11)
Using these relations, we are ready to present a new formulation of ((M+1)-Cyclic):
$\max\limits_{\boldsymbol{x},\boldsymbol{\rho},\boldsymbol{\gamma},\boldsymbol{\Gamma}}\quad$ $\frac{1}{M+1}\sum_{t\in[M+1]}\sum_{i\in[N]}r_{i}u_{i}\gamma_{i}^{t}$
$\operatorname{s.t.}\quad$ $x_{i}^{t}+\sum_{m\in[M]}x^{\tau(m\mid t)}_{i}\leq 1\text{ and }\boldsymbol{x}^{t}\in\mathcal{X}\cap\{0,1\}^{N}$ $\operatorname{for}i\in[N]\text{ and }t\in[M+1]$ (12a)
$\rho^{t}+\sum_{i\in[N]}u_{i}\gamma^{t}_{i}=1\text{ and }0\leq\gamma^{t}_{i}\leq\rho^{t}\quad$ $\operatorname{for}i\in[N]\text{ and }t\in[M+1]$ (12b)
$x^{t}_{i}=\gamma^{t}_{i}+\sum_{j\in[N]}u_{j}\Gamma^{t}_{ij}$ $\operatorname{for}i\in[N]\text{ and }t\in[M+1]$ (12c)
$\max\{0,\gamma^{t}_{i}+\gamma^{t}_{j}-\rho^{t}\}\leq\Gamma^{t}_{ij}\leq\min\{\gamma^{t}_{i},\gamma^{t}_{j}\}\quad$ $\operatorname{for}i\in[N]\,,j\in[N]\setminus\{i\},\text{ and }t\in[M+1]$ (12d)
$\Gamma^{t}_{ii}=\gamma^{t}_{i}$ $\operatorname{for}i\in[N]\text{ and }t\in[M+1].$ (12e)
@@
The constraint (12b) is obtained from relaxing the constraints from (11). The constraint (12c) is derived from the following equality, with the variable $\gamma^{t}_{i}$ replacing the first fractional term and a new variable $\Gamma^{t}_{ij}$ replacing each fractional term in the summation,
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x^{t}_{i}=\frac{x^{t}_{i}}{1+\sum_{k\in[N]}u_{k}x^{t}_{k}}+\sum_{j\in[N]}u_{j}\frac{x_{i}^{t}x_{j}^{t}}{1+\sum_{k\in[N]}u_{k}x^{t}_{k}}\qquad\operatorname{for}\,t\in[M+1]}.$
To obtain (12d), we use the no-purchase variable $\rho^{t}$ to scale the following McCormick envelope of $x_{i}^{t}x_{j}^{t}$ (McCormick 1976):
$\max\{0,x^{t}_{i}+x^{t}_{j}-1\}\leq x^{t}_{i}x^{t}_{j}\leq\min\{x_{i}^{t},x_{j}^{t}\}\qquad\operatorname{for}i\in[N],j\in[N]\setminus\{i\}.$
@@
The last constraint (12e) is obtained by observing that the relation $x^{t}_{i}x^{t}_{i}=x^{t}_{i}$ holds for binary variables. Unlike the formulation (Cycle-Conic), this formulation does not rely on bounds of the no-purchase probability variables. More specifically, in the derivation of (Cycle-Conic), we use $\rho^{t}x^{t}_{i}$ as a nonlinear scaling factor, demanding a linearization using bounds of $\rho^{t}$. In the development of linear inequalities in (12), we avoid using a nonlinear scaling factor and, instead, directly use $\rho^{t}$ as a scaling variable to derive (12d).
@@
The following result shows this bound-free formulation is theoretically tighter than the formulation (Cycle-Conic). One of the most important properties of an mixed integer program (MIP) formulation is the strength of its natural continuous relaxation that is obtained by ignoring the integrality constraints. This is important because a tighter continuous relaxation often indicates a faster convergence of the branch-and-bound algorithm on which most commercial MIP solvers are built (Vielma 2015). Since (HAP) has a maximization objective, a tighter formulation has a smaller continuous relaxation objective value. Let $Z_{\textsc{Bound-free}}$ (resp. $Z_{\textsc{Cycle-Conic}}$) be the optimal objective value of the natural continuous relaxation of formulation (12) (resp. (Cycle-Conic)). The two objective values have the following relationship.
@@
$Z_{\textsc{Bound-free}}\leq Z_{\textsc{Cycle-Conic}}$ if the cross-period constraint is non-overlapping.
@@ 5 Case Study
This section presents a case study using transaction data from a corporation cafeteria. We first estimate the history-dependent effect and then use the results to design assortments. Our estimations indicate that customers have a significant but short-term satiation effect. Furthermore, our findings suggest that including history-dependent effects in choice models will lead to greater assortment varieties and higher revenue.
@@ 5.1 Data Description and Estimation
We cooperated with a Chinese company that provides catering services to over $4,000$ corporation cafeterias and were given access to a dataset for one cafeteria that mainly serves workday luncheons for customers who work nearby. The dataset includes $27,144$ transactions for $53$ main dishes, covering $110$ workdays from July to December 2021. The data for each transaction includes the time, dish, and quantity, as well as a card ID if the dish was purchased with a membership card. The dataset also includes information about the price, main ingredient, and flavor of each dish. In what follows, we use the terms dish and product and also the terms menu and assortment interchangeably.
Our dataset indicates that 84.8% of the transactions came from 418 members, which generated 84.5% of the total revenue. Thus, we focus on the members’ transactions because they are frequent visitors and generate most of the revenue. Our dataset also reveals that the probability of a member dining in the cafeteria more than once for two, three, or four consecutive days was 72.5%, 87.7%, and 91.2%, respectively. This suggests that customers’ dish selections could be influenced by past menus.
@@
$U_{i}^{t}=\boldsymbol{\theta}^{T}X_{i}+\theta_{r}r_{i}+\sum_{m\in[M]}x_{i}^{t-m}\beta^{m}+\xi_{i}^{t},$
@@
where $X_{i}$ denotes covariates that include a fixed effect and the main ingredient and flavor of product $i$; $r_{i}\in[0,1]$ is a normalized price; $\boldsymbol{\theta}$ and $\theta_{r}$ are coefficients of $X_{i}$ and price, respectively; $\boldsymbol{\theta}^{T}X_{i}+\theta_{r}r_{i}$ is the base utility $\beta_{i}^{0}$ defined in Section 2; and $\xi_{i}^{t}$s are i.i.d. standard Gumbel random variables. The history-dependent effect is captured by the term $\sum_{m\in[M]}x_{i}^{t-m}\beta^{m}$, in which $M$ is the memory length. Our model is a standard MNL model if $M=0$. Note that we set the same history-dependent coefficient for all products because not all products have consecutive offering records, and therefore, we cannot estimate the product-wise history-dependent effect.
@@
We utilize members’ transactional data to calibrate the MNL choice model with the utility function presented above. We split the transaction data into two sets: data from July to October for training and data from November to December for test. We first use five-fold cross-validation on the first data set to estimate the utility function, and then we optimize weekly menus based on the estimators. Next, we conduct the five-fold cross-validation again, this time on the second data set, to select the value of $M$ with the best out-of-sample performance, and we set this as the “ground truth” model for evaluation, which is a common approach in policy evaluation (Cao et al. 2022). To address the fact that our dataset lacks no-purchase records, we employ the expectation-maximization (EM) approach, a widely used technique to deal with censored data (van Ryzin and Vulcano 2017, Şimşek and Topaloglu 2018), to calibrate our model.
@@
Table 5.1 summarizes the estimation and five-fold validation results for $M\in\{0,1,2,3,4\}$, using the first data set. Column $\theta_{r}$ presents price coefficient estimators, and columns $\beta^{1}$ to $\beta^{4}$ give history-dependent effects. The columns labeled $\chi^{2}$ and Log-likelihood provide the average out-of-sample value in the five-fold cross-validation (normalized by sample size). The percentage values in parentheses indicate the relative improvement of each model with a history-dependent effect compared to the standard MNL model.
[…]
@@
Estimators and cross-validation results under different memory length Estimators Out-of-sample Performance $M$ $\theta_{r}$ $\beta^{1}$ $\beta^{2}$ $\beta^{3}$ $\beta^{4}$ $\chi^{2}$ Log-likelihood 0 $-2.558^{***}$ 0.0860 -1.7796 (0.0767) 1 $-2.550^{***}$ $-0.475^{***}$ 0.0828 ($3.67\%$) -1.7769 ($0.15\%$) (0.0767) (0.0498) 2 $-2.520^{***}$ $-0.493^{***}$ $-0.366^{***}$ 0.0822 ($4.38\%$) -1.7754 ($0.24\%$) (0.0767) (0.0499) (0.0568) 3 $-2.540^{***}$ $-0.492^{***}$ $-0.358^{***}$ $-0.289^{***}$ 0.0819 ($4.71\%$) -1.7749 ($0.27\%$) (0.0768) (0.0500) (0.0568) (0.0708) 4 $-2.540^{***}$ $-0.495^{***}$ $-0.361^{***}$ $-0.286^{***}$ $-0.146^{***}$ 0.0826 ($3.91\%$) -1.7753 ($0.25\%$) (0.0768) (0.0500) (0.0569) (0.0709) (0.0647) Note: The numbers in parentheses under the estimators report the standard errors of the estimations. “-” indicates that a model does not include the corresponding predictor. ${***}p<0.001$. The columns “$\chi^{2}$” and “Log-likelihood” show the average value of the five-fold cross-validation (normalized by sample size) followed by a percentage in parentheses, which is the relative improvement of each model with a history-dependent effect compared to the standard MNL model.
@@
Table 5.1 shows significant negative history-dependent effects, and the scale of the effects decreases as the time interval between the current and historical offerings increases. The estimation results are consistent for different values of $M$. For instance, $\beta^{1}$ is consistently about $-0.49$ for all $M$. This suggests that the attraction value of a product will shrink by about $\exp(-0.49)=61\%$ if it was offered yesterday. In other words, if a product is sold on two consecutive days, we need to lower its price by at least RMB $(0.492/2.540)\times 15\approx 2.9$ when $M=3$ to compensate for the negative history-dependent effect, where we normalize prices by the maximum price RMB $15$ in the estimation. These results demonstrate the importance of considering history-dependent effects in assortment planning.
@@
The average out-of-sample validation results are reported in the last two columns of Table 5.1. They suggest that history-dependent effects can improve prediction power (i.e., lower $\chi^{2}$ and higher log-likelihoods). When $M=3$, the model presents the lowest $\chi^{2}$ and the highest log-likelihood. This result suggests that a small memory length ($M=3$) can capture the history-dependent effect in customers’ choice.